# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v2

@@ Abstract
Verifying the provenance of content is crucial to the function of many organizations, e.g., educational institutions, social media platforms, firms, etc. This problem is becoming increasingly difficult as text generated by Large Language Models (LLMs) becomes almost indistinguishable from human-generated content. In addition, many institutions utilize in-house LLMs and want to ensure that external, non-sanctioned LLMs do not produce content within the institution. In this paper, we answer the following question: Given a piece of text, can we identify whether it was produced by LLM $A$ or $B$ (where $B$ can be a human)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design zero-shot statistical tests to distinguish between (i) the text generated by two different sets of LLMs $\mathcal{A}$ (in-house) and $\mathcal{B}$ (non-sanctioned), and also (ii) LLM-generated and human-generated texts. We prove that the type I and type II errors for our tests decrease exponentially in the text length. In designing our tests, we derive concentration inequalities on the difference between log-perplexity and the average entropy of the string under $A$. Specifically, for a given string, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. We also show that if $B$ generates the text, then, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Lastly, we present preliminary experimental results using open-source LLMs to support our theoretical results. Practically, our work enables guaranteed (with high probability) finding of the origin of harmful or false LLM-generated text for a text of arbitrary size, which can be useful for combating misinformation as well as compliance with emerging AI regulations.
@@ 1 Introduction
LLM and human-generated texts are becoming indistinguishable. This phenomenon has some concerning societal consequences, including the spread of LLM-generated misinformation and LLM-assisted academic cheating. We need reliable detection methods to distinguish between human and AI-generated content. In addition to distinguishing human vs. LLM-generated text, the ability to differentiate between text generated by a language model vs. another language model is also critical because of the following reasons. First, the ability to find the origin of harmful or false LLM-generated content is essential for legal compliance and mitigation purposes (Hacker et al., 2023; Wang et al., 2023). Identifying the source of harmful content allows for responsibility to be assigned in case of non-compliance with regulations. Second, the quality of outputs can vary significantly between models, and as a result, employing the most appropriate model for the specific task is essential to achieve optimal results. In line with operating the most suitable language model, educational organizations are building their LLMs to secure educational integrity and credibility. For example, the University of Michigan has developed U-M GPT and UM Maizey as its generative AI tools to ensure academic integrity, guarantee user data protection, and assure that the shared information does not train the underlying AI models (UMichigan (2023)). Students are allowed only to use the specialized assistant (here, UM Maizey) to do their assignments, not ChatGPT or any other LLM. This requirement highlights the need for reliable tools to detect the text generated by prohibited LLMs.
@@
To evaluate the possibility of detecting LLM-generated text by human experts in linguistics, Casal and Kessler, 2023 designed an experiment to investigate whether linguists can distinguish human and ChatGPT-generated text and reported an identification rate of only $38.9\%$. Since humans, even experts, perform poorly in detecting LLM-generated text, researchers are investing significant efforts in designing automated detection methods to identify signals that are difficult for humans to recognize.
@@
One way to create a detection method is to train classifiers on labeled training data coming from LLM-generated and human-generated classes. OpenAI itself released an AI classifier model. However, they later discontinued operating the model due to its low accuracy rate (OpenAI ). While recent literature on classification methods with higher accuracy rates exists (e.g., Guo et al., 2023), the limitations of this method make its application practically challenging. The first major limitation is the requirement for training a separate (from the source-model) classifier, especially considering the large and growing number of LLMs, the wide variety of topics and writing styles, and the possibility of prompting LLMs to write in different styles. Furthermore, the requirement for collecting a dataset of human and AI generated passages raises concerns, such as privacy, associated with training models on human data. Finally, Liang et al., 2023 note that because detectors are often evaluated on relatively easy datasets, their performance on out-of-domain samples is often abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using commercial detectors.
@@
An alternative approach for detecting LLM-generated content is watermarking. However, watermarking relies on cooperation from the AI company/owner of the LLM. Although using watermarking for AI regulation is imaginable, current regulations cannot force companies to adopt this technology (Nature, 2024). The described limitations motivate the need for models that do not require training on human data or cooperation from the LLM owner. One such method is zero-shot training (see, e.g., the work of Mitchell et al., 2023).
The majority of the zero-shot detection studies (e.g., Mitchell et al., 2023 and Hans et al., 2024) leverage statistical properties of LLM-generated texts to develop statistical tests for identifying whether a human or an LLM generates a finite-length text. The proposed methods rely on heuristics that perform well in practice, but they do not provide statistical guarantees for their detection mechanism’s accuracy level, making them unfit for any purpose where strong evidence is required (e.g., regulatory compliance).
@@
In this study, we answer the following question: Given a piece of text, can we identify whether it was generated by LLM $A$ or $B$ (where $B$ can be a human or another LLM among a set of LLMs)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single string of text with finite length, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We design tests to distinguish between different LLMs. For this purpose, we assume that we have white-box access to the models in the hypothesis test. In particular, we design composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. We also study the case where we do not have white-box access to all models in the hypothesis set (for example, a human wrote the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
@@
Finally, our theoretical results include establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables. These concentration bounds have been derived for a sequence of independent and identically distributed (iid) random variables by Zhao 2022. However, large language models generate text sequences dependent on the previously generated tokens. Motivated by the problem requirement, we generalize the results in the literature (e.g., Zhao 2022) by proving an exponential decay concentration inequality to bound the tail probability of the difference between the log-likelihood of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables satisfying a martingale structure. Interestingly, as a byproduct of this paper, we address one of the future research directions mentioned by Zhao 2022, namely deriving entropy concentration bounds for martingales.
This paper is organized as follows: In Section 2, we present our mathematical framework and define the critical random variables necessary for deriving our theoretical results. Section 3 consists of two subsections. In Subsection 3.1, we provide concentration bounds when the generator and evaluator models are the same, and in Subsection 3.2, we provide concentration bounds when the generator and evaluator models are different. In Section 4, we present our statistical tests and derive the upper bound on their type I and type II errors. In Section 5, we show the results of our preliminary experiments on an open-source language model. We conclude the paper in Section 6. Appendix A presents the detailed proofs, and Appendix B provides a literature review.
@@ 2.1 Model
Remark 1.
Equation (1) is an application of the Bayes’ rule and holds for any generative model (regardless of whether tokens $Y_{n}$ are sequentially generated). While equation (1) holds for all generative models, because conditional distributions $p_{n}(y)$ are in general not easily accessible, we apply the rule for sequential models. We follow the literature on white-box detection, and assume that we have complete knowledge of the probability law $p^{M}_{n}(\textbf{Y}_{n})$ for any given sequence $\textbf{Y}_{n}$. See, for example, Mitchell et al. 2023; Gehrmann et al. 2019.
@@ 2.2 Background definitions
The perplexity $p^{A}(Y_{N})$ of a (finite length) text string $\textbf{Y}_{N}=[Y_{1},Y_{2},\ldots,Y_{N}]$ with respect to an evaluator generative model $A$ is defined as the per-token inverse likelihood of the string Y. Formally, perplexity with respect to model $A$ is
@@
We aim to design statistical tests to reliably evaluate whether text is generated by model $A$ or by a different model $B$, which can be a different generative model or human. Suppose model $B$ generates a string $\mathbf{Y}$. The cross-entropy of model $B$ and evaluator model $A$ over sub-string $\mathbf{Y}_{N}$ equals
$h_{N}(B,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$ (3)
Note that if model $A$, which is also the evaluator model, generates the string, then cross-entropy and entropy are equivalent and defined as
$h_{N}(A,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}p^{A}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$

### v3

@@ Abstract
Verifying the provenance of content is crucial to the function of many organizations, e.g., educational institutions, social media platforms, firms, etc. This problem is becoming increasingly challenging as text generated by Large Language Models (LLMs) becomes almost indistinguishable from human-generated content. In addition, many institutions utilize in-house LLMs and want to ensure that external, non-sanctioned LLMs do not produce content within the institution. In this paper, we answer the following question: Given a piece of text, can we identify whether it was produced by LLM $A$ or $B$ (where $B$ can be a human)? We model LLM-generated text as a sequential stochastic process with complete dependence on history and design zero-shot statistical tests to distinguish between (i) the text generated by two different sets of LLMs $\mathcal{A}$ (in-house) and $\mathcal{B}$ (non-sanctioned), and also (ii) LLM-generated and human-generated texts. We prove that our tests’ type I and type II errors decrease exponentially as text length increases. In designing our tests, we derive concentration inequalities on the difference between log-perplexity and the average entropy of the string evaluated by $A$. Specifically, for a given string, we demonstrate that if the string is generated by the evaluator model $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in the string length. We also show that if $B$ generates the text, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$. Lastly, we present two sets of experiments: first, we present experiments using open-source LLMs to support our theoretical results, and then we provide experiments in a black-box setting with adversarial attacks. Practically, our work enables guaranteed (with high probability) finding of the origin of harmful or false LLM-generated text for a text of arbitrary size, which can be useful for combating misinformation as well as compliance with emerging AI regulations.
@@ 1 Introduction
LLM text generation tools such as GPT-4 (OpenAI 2023), Llama (Touvron et al. 2023), Gemini (Team et al. 2024) and Mixtral (Jiang et al. 2024) are being widely employed to produce textual content in various domains including news agencies (Newsguard 2024) and academia (Originality.ai 2024). As LLM content generation tools improve, accurately telling whether a text is human-generated or LLM-generated becomes increasingly challenging.
But why do we need to know whether a person or an LLM generated the text? First, LLMs allow people to deliver content they did not produce. In educational settings, for example, this disrupts current accreditation systems. It can also impact the quality of academic peer reviews (Reza et al. 2024). Second, LLMs let people proliferate content. That means an individual can produce significantly more content, including content deliberately tailored to specific audiences, influencing public discussions, e.g., manipulating financial markets (News 2024), political discussions (Times 2023), or consumers’ sentiments (Jakesch et al. (2023)). As stated, not being able to distinguish between LLM and human-generated text yields societal consequences, including the spread of LLM-generated misinformation (Chen and Shu 2024) and LLM-assisted academic cheating Cotton et al. 2024. Therefore, from the policy maker’s perspective, providing methods that reliably distinguish between human and LLM-generated content is crucial. Alongside policymakers, generative AI providers also have incentives to be able to detect AI and human-generated content. That is because retraining on AI-generated content can lead to a phenomenon named "model collapse," first reported by Shumailov et al., 2023. To avoid model collapse, AI providers need to exclude AI-generated content from their training sets, requiring them to identify such content reliably.
@@
In addition to distinguishing human vs. LLM-generated text, henceforth referred to as the detection problem, the ability to differentiate between text generated by a language model vs. another language model, henceforth referred to as the attribution problem, is also critical for the following reasons. First, finding the origin of harmful or false LLM-generated content is essential for legal compliance and mitigation purposes (Hacker et al., 2023; Wang et al., 2023a). Identifying the source of harmful content allows for responsibility to be assigned in case of non-compliance with regulations. Second, the quality of outputs can vary significantly between models, and as a result, employing the most appropriate model for the specific task is essential to achieving optimal results. In line with operating the most suitable language model, educational organizations are building their LLMs to secure educational integrity and credibility. For example, the University of Michigan has developed U-M GPT and UM Maizey as its generative AI tools to ensure academic integrity, guarantee user data protection, and ensure that the shared information does not train the underlying AI models (UMichigan 2023). Students can use only the specialized assistant (here, UM Maizey) to do their assignments, not ChatGPT or any other LLM. This requirement highlights the need for reliable tools to detect the text generated by prohibited LLMs. Again, alongside policymakers, generative AI providers have incentives to be able to detect the text generated by their own language model to gauge the uptake of their systems, which is a commercially important measure of performance.
Figure 1: (Left) LLM-generated Text Detection focuses on detecting whether a text has been generated by an LLM or by a human; (Right) LLM-generated Text Attribution focuses on detecting the LLM source that generated the text among candidate sources.
@@
To evaluate the possibility of detecting LLM-generated text by human linguistics experts, Casal and Kessler, 2023 designed an experiment to investigate whether linguists can distinguish human and ChatGPT-generated text and reported an identification rate of only $38.9\%$. Since humans, even experts, perform poorly in detecting LLM-generated text, researchers are investing significant efforts in designing automated detection methods to identify signals that are difficult for humans to recognize.
@@
One approach to creating a detection method is to train classifiers on labeled training data from LLM-generated and human-generated classes (Bakhtin et al. 2019; Jawahar et al. 2020; Uchendu et al. 2020; Fagni et al. 2021; Sadasivan et al. 2023; Guo et al. 2023; Verma et al. 2023). The limitations of supervised detection approaches make their application practically challenging. First is the requirement for training a separate (from the source-model) classifier, especially considering the large and growing number of LLMs, the wide variety of topics and writing styles, and the possibility of prompting LLMs to write in different styles. Furthermore, the requirement for collecting a dataset of human and AI-generated passages raises concerns, such as privacy, associated with training models on human data. Finally, Liang et al., 2023 notes that because detectors are often evaluated on relatively easy datasets, their performance on out-of-domain samples is often abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using commercial detectors.
@@
An alternative detection method is watermarking (Kirchenbauer et al. 2023), in which an LLM embeds hidden signals into the text during generation, and the signals are identifiable by an algorithm while indistinguishable to humans. Watermarking relies on cooperation from the AI company/owner of the LLM, and current regulations cannot force companies to adopt this technology (Nature, 2024). The described limitations motivate the need for models that do not require training on human data or cooperation from the LLM owner. One such method is zero-shot training, which does not require additional data collection (see, e.g., the work of Mitchell et al., 2023).
The majority of zero-shot detection studies leverage statistical properties of LLM-generated text for identification, including likelihood curvature (Mitchell et al. 2023), log-likelihood (Solaiman et al. 2019), rank-likelihood (Gehrmann et al. 2019), log-likelihood ratio ranking (Su et al. 2023), entropy and Kullback-Leibler (KL) divergence (Lavergne et al. 2008), and perplexity (Vasilatos et al. 2023; Wang et al. 2023b). While most methods focus on a single statistical property, Hans et al. (2024) introduces a normalized measure of perplexity by dividing log-perplexity by average cross-entropy. The rationale is that LLM-generated text is predictable to the LLM (low perplexity), whereas human text is more surprising (higher perplexity). However, high perplexity as a human-authorship signal is unreliable due to prompt dependency, as illustrated by the "Capybara Problem" (Hans et al. 2024). For instance, given the prompt "Can you write about a capybara astrophysicist?", an LLM may generate "The capybara studied dark matter in Andromeda." Without the prompt, the words "capybara" and "astrophysicist" seem unexpectedly high in perplexity, falsely suggesting human authorship. While existing methods, perform well empirically, they remain heuristic-based and lack formal guarantees.
To address this limitation, our study provides a theoretical foundation for distinguishing between LLM-generated and human-written text. In our study, we introduce a framework for determining whether a given string was produced by LLM $A$ or another source $B$ (which may be a human or another LLM). Specifically, for a given string, we demonstrate that if the string is generated by $A$, the log-perplexity of the string under $A$ converges to the average entropy of the string under $A$, except with an exponentially small probability in string length. We also show that if $B$ generates the text, except with an exponentially small probability in string length, the log-perplexity of the string under $A$ converges to the average cross-entropy of $B$ and $A$.
@@
We model LLM-generated text as a sequential stochastic process with complete dependence on history and design statistical tests that take a single string of text with finite length, a prompt, and a given LLM as input and assess whether the given LLM produced the text. We design tests to distinguish between different LLMs. For this purpose, we assume that we have white-box access to the models in the hypothesis test. In particular, we design composite tests that determine whether a text is generated by a model that belongs to a set of models $\mathcal{A}$ or a model that belongs to a disjoint set of models $\mathcal{B}$. We also study the case where we do not have white-box access to all models in the hypothesis set (for example, a human wrote the text) and design a composite statistical test to identify whether the text is generated by a model $A$ or not.
@@
Finally, our theoretical results include establishing concentration bounds for the difference between the log-likelihood of a sequence of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables. These concentration bounds have been derived for a sequence of independent and identically distributed (iid) random variables by Zhao 2022. However, large language models generate text sequences dependent on the previously generated tokens. Motivated by the problem requirement, we generalize the results in the literature (e.g., Zhao 2022) by proving an exponential decay concentration inequality to bound the tail probability of the difference between the log-likelihood of discrete random variables on a finite alphabet and the negative entropy for non-independent random variables satisfying a martingale structure. Interestingly, as a byproduct of this paper, we address one of the future research directions mentioned by Zhao 2022.
Mudireddy et al. (2024) shows that the log-perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions when the evaluator and the generator model are the same. They note that their result relies on sufficiently large text samples to reach asymptotic behavior. This version of their manuscript, with correct proofs, was made online at the same time/after the initial release of our manuscript. We generalize this work to the case with an arbitrary length text and the case where generator and evaluator models are different by showing the convergence of log-perplexity to cross-entropy between the generator and the evaluator models, except with an exponentially small probability in the string length.
This paper is organized as follows: In Section 2, we present our mathematical framework and define the critical random variables necessary for deriving our theoretical results. Section 3 consists of two subsections. In Subsection 3.1, we provide concentration bounds when the generator and evaluator models are the same, and in Subsection 3.2, we provide concentration bounds when the generator and evaluator models are different. In Section 4, we present our statistical tests and derive the upper bound on their type I and type II errors. In Section 5, we show the results of our experiments on open-source language models. We conclude the paper in Section 6. Appendix A presents the detailed proofs, and Appendix B provides a literature review.
@@ 2.1 Model
Note that Equation (1) is an application of the Bayes’ rule and holds for any generative model (regardless of whether tokens $Y_{n}$ are sequentially generated). While equation (1) holds for all generative models, because conditional distributions $p_{n}(y)$ are in general not easily accessible, we apply the rule for sequential models. We follow the literature on white-box detection, and assume that we have complete knowledge of the probability law $p^{M}_{n}(\textbf{Y}_{n})$ for any given sequence $\textbf{Y}_{n}$. See, for example, Mitchell et al. 2023; Gehrmann et al. 2019.
@@ 2.2 Background definitions
We aim to design statistical tests to reliably evaluate whether text is generated by model $A$ or by a different model $B$, which can be a different generative model or human.
Cross-entropy. Suppose model $B$ generates a string $\mathbf{Y}$. The cross-entropy of model $B$ and evaluator model $A$ over sub-string $\mathbf{Y}_{N}$ equals
$h_{N}(B,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{n}\in\mathcal{X}}p^{B}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$ (2)
Intuitively, cross-entropy measures how surprising the token predictions of $A$ are when observed by $B$.
Entropy. Note that if model $A$, is both the generator and the evaluator of the string, then cross-entropy and entropy are equivalent and defined as
$h_{N}(A,A)(\textbf{Y}_{N})=-\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{n}\in\mathcal{X}}p^{A}_{n}(y_{n})\log(p^{A}_{n}(y_{n})).$
Perplexity. The perplexity $p^{A}(Y_{N})$ of a (finite length) text string $\textbf{Y}_{N}=[Y_{1},Y_{2},\ldots,Y_{N}]$ with respect to an evaluator model $A$ is defined as the per-token inverse likelihood of the string Y. Formally, perplexity with respect to model $A$ is